OpenStudy (anonymous):

Your piggy bank has a total of 18 coins in it; some are dimes and some are nickels. If you have a total of $1.40, how many nickels do you have? 14 12 10 8

OpenStudy (austinl):

1.4=.1x+.05y

OpenStudy (anonymous):

10 dimes 8 nickels

OpenStudy (anonymous):

Let x=number of nickels. Since there are 18 coins total, then 18-x=number of dimes. Since we know the total value is 140 cents, then 5 times the number of nickels plus 10 times the number of dimes equals 140. \[5x+10(18-x)=140\] Distribute and simplify: \[5x+180-10x=140\] \[-5x=-40\] \[x=\]

OpenStudy (anonymous):

I got a decimal

OpenStudy (anonymous):

I don't know why this didn't post: \[x=-40/-5=8 nickels\]

OpenStudy (anonymous):

okay thank you!